# Unsupervised Document and Template Clustering using Multimodal Embeddings

### v1

@@ 3.3.2 DBSCAN
DBSCAN (Density-Based Spatial Clustering of Applications with Noise) is a density-based clustering algorithm that groups together data points that are closely packed together, marking as outliers points that lie alone in low-density regions. DBSCAN defines clusters based on the density of data points in the feature space. It relies on two parameters:
@@ 4 Results and Discussion
This section comprises four parts. First, we formalize the external and internal evaluation metrics that will be used throughout (§4.1). Section 4.2 then presents the template-level clustering experiments, whose objective is to group pages that instantiate an identical layout within the same document category. The datasets specific to this task are introduced at the beginning of this subsection. Section 4.3 addresses document-level clustering, where each single-page document is clustered according to its general type, and the corresponding corpora are likewise detailed in situ. Finally, Section 4.4 reports ablation studies that isolate architectural capacity and pooling strategy: we contrast the base and large-scale LayoutLMv1 checkpoints and compare the mean pooling of the final hidden states with the conventional [CLS] token.
The OCR model used to extract textual content from the document images is the CRNN model [38], chosen for its robustness across various document layouts and its efficiency. Apart from a light deskew applied only when obvious rotation was detected, no preprocessing, normalization or denoising is performed. Note that $k$-Means is supplied with the ground-truth number of clusters, giving it a distinct advantage over DBSCAN. In all experiments $k$-Means is initialized with the greedy $k$-Means++ scheme.
@@ 4.1 Evaluation Metrics
Furthermore, we use the homogeneity score (HS) and the completeness score (CS), both introduced by Rosenberg and Hirschberg [42]. HS assesses the extent to which each cluster contains only members of a single class, while CS assesses whether all members of a given class are assigned to the same cluster. These two metrics complement each other, and their harmonic mean is captured by the V-measure (VM), providing a balanced evaluation of both cluster homogeneity and completeness.
@@ 4.2.2 Experimental results
The three corpora enable a controlled inspection of how text-only, layout-aware, vision-only, and vision-language representations shape the topology of template-level embeddings, and consequently, the behavior of $k$-Means versus DBSCAN.
@@
Table 3: Template-level clustering results on the clean FATURA synthetic-invoice corpus. For every model + algorithm pair we report adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), V-measure (VM), silhouette score (SS), the number of predicted clusters (PC) – ground truth having 50 clusters – and the percentage of noise points returned by DBSCAN. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS VM SS PC % Noise
SBERT $k$-Means 0.7615 0.9306 0.9180 0.9436 0.9306 0.2584
SBERT DBSCAN 0.0119 0.2529 0.1520 0.7527 0.2529 0.0286 9 6.90
LayoutLMv1 $k$-Means 0.0665 0.4386 0.9612 0.2841 0.4386 0.1147
LayoutLMv1 DBSCAN 0.1380 0.7029 0.5741 0.9063 0.7029 0.0124 30 15.92
LayoutLMv3 $k$-Means 0.7422 0.9091 0.9010 0.9174 0.9091 0.6007
LayoutLMv3 DBSCAN 0.8579 0.9479 0.9960 0.9042 0.9479 0.6811 94 0.59
DiT $k$-Means 0.9644 0.9882 0.9872 0.9893 0.9882 0.6503
DiT DBSCAN 0.9179 0.9667 1.0000 0.9356 0.9667 0.7375 79 0
Donut $k$-Means 0.9708 0.9921 0.9907 0.9935 0.9921 0.6937
Donut DBSCAN 0.9240 0.9856 0.9717 1.0000 0.9856 0.6756 46 0
ColPali $k$-Means 0.7862 0.9241 0.9151 0.9334 0.9241 0.1698
ColPali DBSCAN 0.0050 0.0788 0.0437 0.3984 0.0788 0.1176 2 15.38
@@
Table 3 reports the complete set of template-level clustering scores obtained on the clean FATURA corpus. It shows that the raster signal coincides almost perfectly with the template identity, and the purely visual encoders are virtually indistinguishable from a supervised classifier. DiT and Donut achieve ARI = 0.9644 / 0.9708, NMI = 0.9882 / 0.99921 and V measure 0.9882 / 0.972 with $k$-Means. When DBSCAN must infer the number of clusters, these models still exceed 0.90 ARI and predict 46–53 clusters while declaring less than 1% noise, indicating that visual features carve well-separated tight basins of density. LayoutLMv3, which fuses lexical, positional, and visual cues, follows at ARI = 0.7422. Although the text-only SBERT retains a respectable 0.7615 ARI, its DBSCAN run collapses to 9 clusters, ARI = 0.0119, and 6.9% noise. ColPali, a vision-language retriever that should, in principle, exploit the raster image, behaves unexpectedly as SBERT: its $k$-Means ARI is 0.7862 and its DBSCAN ARI drops below 0.02.
Table 4: Template-level clustering on Transformed FATURA, the FATURA invoices after heavy photometric and geometric augmentation. Column definitions are identical to Table 3. Ground truth has 50 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS VM SS PC % Noise
SBERT $k$-Means 0.7378 0.8642 0.8571 0.8715 0.8642 0.1437
SBERT DBSCAN 0.0043 0.0791 0.0446 0.3464 0.0791 0.1183 4 13.85
LayoutLMv1 $k$-Means 0.3574 0.6527 0.6443 0.6613 0.6527 0.0694
LayoutLMv1 DBSCAN 0.0030 0.0435 0.0243 0.2114 0.0435 0.2648 3 15.63
LayoutLMv3 $k$-Means 0.4333 0.6735 0.6658 0.6814 0.6735 0.2411
LayoutLMv3 DBSCAN 0.0986 0.4862 0.3809 0.6718 0.4862 0.1360 24 14.52
DiT $k$-Means 0.2967 0.5261 0.5185 0.5340 0.5261 0.1759
DiT DBSCAN 0.0006 0.0166 0.0091 0.0947 0.0166 0.2995 4 11.07
Donut $k$-Means 0.1556 0.3596 0.3529 0.3666 0.3596 0.1986
Donut DBSCAN 0.0004 0.0073 0.0040 0.0380 0.0073 0.3060 2 14.45
ColPali $k$-Means 0.7176 0.8672 0.8587 0.8759 0.8672 0.1242
ColPali DBSCAN 0.0014 0.0357 0.0193 0.2289 0.0357 0.1818 2 10.25
@@
The ranking reverses on Transformed FATURA, where folds, scanner shadows and speckle blur template boundaries without altering token identities. Table 4 presents the clustering results in Transformed FATURA, illustrating how heavy visual perturbations degrade or reshape performance relative to the clean baseline. SBERT’s $k$-Means ARI decreases by only three points to 0.7378 and its NMI remains 0.8642, confirming lexical robustness. ColPali exhibits the same stability, suggesting the hypothesis that its embeddings are effectively text-dominated in this setting. In contrast, DiT, Donut and LayoutLMv3 lose 0.45–0.80 ARI and their completeness plummets, while homogeneity remains moderate: visual corruption fragments each template into many pure but incomplete islands. DBSCAN now predicts between 2 and 24 clusters for every encoder. In addition, it never exceeds 0.099 ARI, and labels up to 15.6% of pages as noise. The large gap between homogeneity and completeness is therefore an immediate diagnostic of modality failure: high homogeneity but low completeness signals that vision-centric models over-segment when the raster channel is unreliable.
@@
Table 5: Template-level clustering on the real scanned Mexican death certificates corpus. Metrics and column conventions follow Tables 34. Ground truth has 9 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS VM SS PC % Noise
SBERT $k$-Means 0.2210 0.5297 0.7893 0.3986 0.5297 0.0926
SBERT DBSCAN 0.7462 0.6810 0.7311 0.6374 0.6810 0.2240 8 14.00
LayoutLMv1 $k$-Means 0.1533 0.3972 0.5713 0.3044 0.3972 0.3669
LayoutLMv1 DBSCAN 0.0191 0.1464 0.1393 0.1541 0.1464 0.5422 5 6.40
LayoutLMv3 $k$-Means 0.2858 0.5637 0.8136 0.4313 0.5637 0.2996
LayoutLMv3 DBSCAN 0.7713 0.6995 0.7999 0.6216 0.6995 0.4493 11 9.60
DiT $k$-Means 0.0409 0.1796 0.2825 0.1317 0.1796 0.2011
DiT DBSCAN -0.0386 0.0513 0.0346 0.0992 0.0513 0.2772 2 11.20
Donut $k$-Means 0.1883 0.4793 0.7267 0.3576 0.4793 0.3238
Donut DBSCAN 0.9403 0.8668 0.9269 0.8139 0.8668 0.2532 11 4.00
ColPali $k$-Means 0.2083 0.5300 0.8020 0.3958 0.5300 0.1215
ColPali DBSCAN 0.6771 0.6218 0.5364 0.7396 0.6218 0.3440 4 6.40
@@
Table 5 lists the clustering results in the Mexican death certificate dataset. The scanned death certificates corpus, whose legal boilerplate renders lexical content nearly invariant across forms, restores the advantage to models that encode spatial and visual structure. The textual clusters of SBERT are coarse ($k$-Means ARI = 0.221; completeness = 0.3986). LayoutLMv3 exploits absolute coordinates to improve to ARI = 0.2858 under oracle k and, crucially, to ARI = 0.7713 with DBSCAN while discovering 11 clusters (ground truth = 9) and flagging less than 10% outliers. Donut attains the best overall performance: DBSCAN yields ARI = 0.9403, NMI = 0.8668, homogeneity = 0.9269, completeness = 0.8139, only 4% noise, and 11 clusters that align one-to-one with human labels except for a benign over-split of two forms. Pure-image DiT, lacking lexical anchoring, collapses to two clusters with $-0.039$ ARI and 11% noise, showing that visual tokens alone cannot resolve subtle scanner artefacts. ColPali remains below the other vision-aware encoders (DBSCAN ARI = 0.312, V-measure = 0.6218), but still better than LayoutLMv1.
@@
Methodologically, the oracle $k$-Means inflates absolute performance and masks the modality weaknesses that surface as soon as k must be inferred. DiT and Donut lose more than 0.45 ARI in Transformed FATURA when the oracle disappears, while Donut gains 0.75 ARI on Mexican death certificates by switching from $k$-Means to DBSCAN. Joint inspection of homogeneity, completeness and the proportion of noise points therefore offers actionable insight that a single aggregate score would obscure, revealing whether an embedding fragments templates or collapses them into amorphous basins.

### v3

@@ 3.3.2 DBSCAN
DBSCAN (Density-Based Spatial Clustering of Applications with Noise) [11] is a density-based clustering algorithm that groups together data points that are closely packed together, marking as outliers points that lie alone in low-density regions. DBSCAN defines clusters based on the density of data points in the feature space. It relies on two parameters:
@@
3.3.3 HDBSCAN + $k$-NN
HDBSCAN [12] extends DBSCAN by building a hierarchy of clusters over a range of density thresholds and then extracting a flat clustering based on cluster stability. In practice, HDBSCAN may return a substantial number of points labeled as noise. To mitigate this, we adopt a two-stage procedure that combines HDBSCAN with a post hoc assignment of $k$-nearest neighbors ($k$-NN) [13].
  1. 1.
    Run HDBSCAN with a chosen metric and hyperparameters (e.g., min_cluster_size, min_samples) to obtain initial cluster labels, where noise points are assigned label 1-1.
  2. 2.
    Train a $k$-NN classifier on the non-noise points (labels $\neq-1$), using their HDBSCAN labels as targets.
  3. 3.
    Predict cluster labels for all points labeled as noise by HDBSCAN using the trained $k$-NN model.
This hybrid approach preserves HDBSCAN’s ability to discover clusters without specifying $k$ while ensuring that, after the $k$-NN reassignment, no data points remain unlabeled as noise. In settings where HDBSCAN provides soft cluster membership probabilities, these can optionally be used as class weights during $k$-NN training.
3.3.4 BIRCH
BIRCH (Balanced Iterative Reducing and Clustering using Hierarchies) [14] incrementally builds a compact summary of the dataset called a Clustering Feature (CF) tree. Each node stores sufficient statistics (number of points, linear sum, and squared sum) that enable efficient updates and distance computations.
  1. 1.
    CF-tree construction: Insert data points one by one into the CF tree using a specified distance threshold $T$ and branching factor $B$, merging or splitting leaf entries to maintain compact summaries.
  2. 2.
    Global clustering (optional): Apply a global clustering algorithm (e.g., $k$-Means) to the leaf entries (subcluster centroids) to obtain the final partition; alternatively, the leaf entries themselves can serve as the final clustering.
BIRCH is well-suited to large or streaming datasets due to its single-pass, memory-efficient construction and its ability to reduce data while preserving cluster structure. Its limitations include a preference for roughly spherical clusters and sensitivity to the threshold parameter $T$, which controls leaf entry compactness.
@@ 4 Results and Discussion
This section comprises six parts. First, we describe the procedure adopted for hyperparameter selection (§4.1), detailing how the clustering algorithms were tuned and, in the case of $k$-Means, how initialization was handled. We then formalize the external and internal evaluation metrics that will be used throughout (§4.2). Section 4.3 presents the template-level clustering experiments, whose objective is to group pages that instantiate an identical layout within the same document category; the datasets specific to this task are introduced at the beginning of this subsection. Section 4.4 addresses document-level clustering, where each single-page document is clustered according to its general type, and the corresponding corpora are likewise detailed in situ. Section 4.5 reports ablation studies that isolate architectural capacity and pooling strategy: we contrast the base and large-scale LayoutLMv1 checkpoints and compare the mean pooling of the final hidden states with the conventional [CLS] token. Finally, Section 4.6 discusses generalization beyond single-page templates, positioning with respect to weakly and semi-supervised alternatives, post-clustering consolidation, and robustness to OCR noise and multilingual inputs.
The OCR model used to extract textual content from the document images is the CRNN model [49], chosen for its robustness across various document layouts and its efficiency. Apart from a light deskew applied only when obvious rotation was detected, no preprocessing, normalization or denoising is performed. Note that $k$-Means and BIRCH are supplied with the ground-truth number of clusters, giving it a distinct advantage over DBSCAN and HDBSCAN + $k$-NN. In all experiments $k$-Means is initialized with the greedy $k$-Means++ scheme.
4.1 Hyperparameter Selection
For clustering algorithms that do not require the number of clusters as input, a robust hyperparameter selection process was essential to ensure optimal performance. To this end, a systematic grid search was implemented to tune the relevant parameters, with the objective of maximizing the internal validation metric, namely the silhouette score. This ensured that the parameterization adhered to a strictly unsupervised methodology and reflected the inherent structure of the data. For DBSCAN, the grid search explored a range of values for both $\epsilon$ and minPts, while for HDBSCAN the search concentrated on different values of the min_cluster_size and min_samples parameters, which largely governs the resulting cluster structure. In the case of BIRCH, a grid search was likewise carried out on the threshold parameter, which controls the compactness of the subclusters in the CF-tree. However, unlike DBSCAN and HDBSCAN, the final evaluation of BIRCH was performed in an oracle mode, in which the ground-truth number of clusters was supplied during the global clustering step. This evaluation setup was aligned with that of $k$-Means, which was also assessed exclusively under an oracle condition. In all experiments, $k$-Means was initialized with the greedy $k$-Means++ scheme to improve convergence and stability.
@@ 4.2 Evaluation Metrics
Furthermore, we use the homogeneity score (HS) and the completeness score (CS), both introduced by Rosenberg and Hirschberg [53]. HS assesses the extent to which each cluster contains only members of a single class, while CS assesses whether all members of a given class are assigned to the same cluster.
@@ 4.3.2 Experimental Results
The three corpora enable a controlled inspection of how text-only, layout-aware, vision-only, and vision-language representations shape the topology of template-level embeddings, and consequently, the behavior of each clustering algorithm.
@@
Table 3: Template-level clustering results on the clean FATURA synthetic-invoice corpus. Metrics are adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), silhouette score ($S$), the number of predicted clusters (PC) and the percentage of noise points returned by DBSCAN. Ground truth has 50 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS $S$ PC % Noise
SBERT $k$-Means 0.7615 0.9306 0.918 0.9436 0.2584
SBERT DBSCAN 0.0119 0.2529 0.152 0.7527 0.0286 9 6.9
SBERT HDBSCAN + $k$-NN 0.8634 0.9490 0.9774 0.9223 0.2646 75
SBERT BIRCH 0.8551 0.9462 0.9418 0.9507 0.2719
LayoutLMv1 $k$-Means 0.8576 0.9604 0.9499 0.9710 0.2763
LayoutLMv1 DBSCAN 0.1380 0.7029 0.5741 0.9063 0.0124 30 15.92
LayoutLMv1 HDBSCAN + $k$-NN 0.8829 0.9703 0.9601 0.9806 0.2932 49
LayoutLMv1 BIRCH 0.9996 0.9997 0.9997 0.9997 0.2983
LayoutLMv3 $k$-Means 0.7422 0.9091 0.9010 0.9174 0.6007
LayoutLMv3 DBSCAN 0.8579 0.9479 0.9960 0.9042 0.6811 94 0.59
LayoutLMv3 HDBSCAN + $k$-NN 0.8334 0.9429 0.9998 0.8921 0.6757 104
LayoutLMv3 BIRCH 0.7563 0.9224 0.9123 0.9328 0.5479
DiT $k$-Means 0.9644 0.9882 0.9872 0.9893 0.6503
DiT DBSCAN 0.9179 0.9667 1 0.9356 0.7375 79 0
DiT HDBSCAN + $k$-NN 0.7367 0.9052 1 0.8269 0.7305 164
DiT BIRCH 0.9923 0.9973 0.9971 0.9974 0.6608
Donut $k$-Means 0.9708 0.9921 0.9907 0.9935 0.6937
Donut DBSCAN 0.9240 0.9856 0.9717 1 0.6756 46 0
Donut HDBSCAN + $k$-NN 0.9385 0.9742 1 0.9496 0.6860 75
Donut BIRCH 0.9759 0.9950 0.9929 0.9970 0.6887
ColPali $k$-Means 0.7862 0.9241 0.9151 0.9334 0.1698
ColPali DBSCAN 0.0050 0.0788 0.0437 0.3984 0.1176 2 15.38
ColPali HDBSCAN + $k$-NN 0.3696 0.7202 0.6616 0.7902 0.1283 35
ColPali BIRCH 0.7735 0.9004 0.8945 0.9062 0.1744
Gemma3 $k$-Means 0.8496 0.9617 0.9527 0.9708 0.4234
Gemma3 DBSCAN 0.9388 0.9718 0.9886 0.9556 0.4331 71 2.14
Gemma3 HDBSCAN + $k$-NN 0.9812 0.9924 1 0.9849 0.4657 56
Gemma3 BIRCH 1 1 1 1 0.4707
InternVL3 $k$-Means 0.9261 0.9790 0.9756 0.9824 0.5308
InternVL3 DBSCAN 0.9310 0.9786 0.9922 0.9653 0.5502 61 0.24
InternVL3 HDBSCAN + $k$-NN 0.9510 0.9826 1 0.9659 0.5526 62
InternVL3 BIRCH 0.9518 0.9897 0.9858 0.9937 0.5533
FATURA.
@@
Table 3 reports the complete set of template-level clustering scores obtained on the clean FATURA corpus. It shows that the raster signal coincides almost perfectly with the template identity, and the purely visual encoders are virtually indistinguishable from a supervised classifier. ARIs are frequently very high, and multiple combinations have a near-perfect result or are perfect. The standout is Gemma3 + BIRCH, which achieves perfect agreement across all external metrics, with $k=50$ provided. DiT and Donut achieve ARI = 0.9923 / 0.9759, NMI = 0.9973 / 0.995, HS = 0.9971 / 0.9929, and CS = 0.9974 / 0.997 with BIRCH, which is superior to $k$-Means. When DBSCAN must infer the number of clusters, these models still exceed 0.90 ARI and predict 46–79 clusters while declaring 0% noise, indicating that visual features form compact, well-separated regions in the embedding space. LayoutLMv3, which fuses lexical, positional, and visual cues, follows at ARI = 0.7422. SBERT performs decently but behind multimodal models: best ARI = 0.8634 with HDBSCAN + $k$-NN, while $k$-Means lags at 0.7615 ARI. Its DBSCAN run collapses to 9 clusters, ARI = 0.0119, and 6.9% noise. ColPali, a vision-language retriever that should, in principle, exploit the raster image, behaves unexpectedly as SBERT: its $k$-Means ARI is 0.7862 and its DBSCAN ARI drops below 0.02. When the value of $k$ must be inferred, HDBSCAN combined with $k$-NN significantly outperforms DBSCAN, achieving a homogeneity score of 0.6616 and a completeness score of 0.7902.
Algorithmically, two patterns stand out. First, BIRCH and $k$-Means, when provided with $k$ = 50, consistently rank among the best performers – frequently achieving ARI scores greater than 0.85 and occasionally reaching perfect scores, as seen with Gemma3. Second, HDBSCAN + $k$-NN almost always improves over plain DBSCAN on ARI and NMI, while driving the percentage of noise points to 0% by design (since $k$-NN assigns outliers back to clusters). Where DBSCAN discards between roughly 0.2% and 2.1% of points as noise for the best performers (e.g., InternVL3 DBSCAN = 0.24%, Gemma3 DBSCAN = 2.14%), HDBSCAN + $k$-NN neutralizes that loss without decreasing ARI or NMI; in fact it often slightly improves them (e.g., SBERT from ARI 0.0119 with DBSCAN to 0.8634 with HDBSCAN + $k$-NN).
Table 4: Template-level clustering on Transformed FATURA. Metrics are adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), silhouette score ($S$), the number of predicted clusters (PC) and the percentage of noise points returned by DBSCAN. Ground truth has 50 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS $S$ PC % Noise
SBERT $k$-Means 0.7378 0.8642 0.8571 0.8715 0.1437
SBERT DBSCAN 0.0043 0.0791 0.0446 0.3464 0.1183 4 13.85
SBERT HDBSCAN + $k$-NN 0 0.0011 0.0006 0.1356 0.5362 2
SBERT BIRCH 0.6878 0.8510 0.8440 0.8581 0.1322
LayoutLMv1 $k$-Means 0.3574 0.6527 0.6443 0.6613 0.0694
LayoutLMv1 DBSCAN 0.0030 0.0435 0.0243 0.2114 0.2648 3 15.63
LayoutLMv1 HDBSCAN + $k$-NN 0 0.0012 0.0006 0.1327 0.9781 2
LayoutLMv1 BIRCH 0.5141 0.7806 0.7729 0.7885 0.0590
LayoutLMv3 $k$-Means 0.4333 0.6735 0.6658 0.6814 0.2411
LayoutLMv3 DBSCAN 0.0986 0.4862 0.3809 0.6718 0.1360 24 14.52
LayoutLMv3 HDBSCAN + $k$-NN 0.2092 0.7094 0.6668 0.7579 0.3418 66
LayoutLMv3 BIRCH 0.4545 0.7205 0.7089 0.7326 0.2289
DiT $k$-Means 0.2967 0.5261 0.5185 0.5340 0.1759
DiT DBSCAN 0.0006 0.0166 0.0091 0.0947 0.2995 4 11.07
DiT HDBSCAN + $k$-NN 0 0.0011 0.0005 0.1333 0.8206 2
DiT BIRCH 0.2849 0.5631 0.5491 0.5778 0.1641
Donut $k$-Means 0.1556 0.3596 0.3529 0.3666 0.1986
Donut DBSCAN 0.0004 0.0073 0.0040 0.0380 0.3060 2 14.45
Donut HDBSCAN + $k$-NN 0 0.0023 0.0012 0.1469 0.7262 3
Donut BIRCH 0.1791 0.4008 0.3866 0.416 0.1412
ColPali $k$-Means 0.7176 0.8672 0.8587 0.8759 0.1242
ColPali DBSCAN 0.0014 0.0357 0.0193 0.2289 0.1818 2 10.25
ColPali HDBSCAN + $k$-NN 0 0.0010 0.0005 0.1905 0.7245 2
ColPali BIRCH 0.7022 0.8448 0.8373 0.8525 0.1155
Gemma3 $k$-Means 0.1521 0.4254 0.4190 0.4319 0.1086
Gemma3 DBSCAN 0.0030 0.0728 0.0397 0.4355 0.2817 3 5.51
Gemma3 HDBSCAN + $k$-NN 0.0031 0.0741 0.0387 0.8909 0.4007 2
Gemma3 BIRCH 0.2295 0.5763 0.5618 0.5916 0.0673
InternVL3 $k$-Means 0.2601 0.5263 0.5197 0.5330 0.1457
InternVL3 DBSCAN 0.0029 0.0748 0.0400 0.5612 0.5561 3 3.16
InternVL3 HDBSCAN + $k$-NN 0 0.0032 0.0016 0.4481 0.5840 2
InternVL3 BIRCH 0.3426 0.6356 0.6251 0.6465 0.1352
Transformed FATURA.
@@
The ranking reverses on Transformed FATURA, where folds, scanner shadows and speckle blur template boundaries without altering token identities. Table 4 presents the clustering results in Transformed FATURA, illustrating how heavy visual perturbations degrade or reshape performance relative to the clean baseline. SBERT’s $k$-Means ARI decreases by only three points to 0.7378 and its NMI remains 0.8642, confirming lexical robustness. ColPali exhibits the same stability, suggesting the hypothesis that its embeddings are effectively text-dominated in this setting. LayoutLMv3 also remains relatively stable with $k$-Means, exhibiting moderate to high NMI but noticeably lower ARI compared to the top two methods. This reflects ongoing fragmentation and merging effects, which ARI is more sensitive to and penalizes more strongly. In contrast, DiT and Donut suffer a sharp decline in all metrics.
DBSCAN now predicts between 2 and 24 clusters for every encoder. In addition, it never exceeds 0.099 ARI, and labels up to 15.6% of pages as noise. The large decrease in completeness and homogeneity is an immediate diagnostic of modality failure: it indicates that vision-centric models under-segment when visual information is unreliable, grouping distinct templates into overly broad clusters. LayoutLMv3 is the only model that consistently maintains moderate to high values across ARI, homogeneity score, and completeness score. InternVL3 performs particularly poorly: its best ARI on Transformed FATURA is just 0.3426 with BIRCH. $k$-Means yields 0.2601, while DBSCAN and HDBSCAN + $k$-NN perform even worse, with ARIs of 0.0029 and 0, respectively.
@@
Table 5: Template-level clustering on the real scanned Mexican death certificates corpus. Metrics are adjusted rand index (ARI), normalized mutual information (NMI), homogeneity score (HS), completeness score (CS), silhouette score ($S$), the number of predicted clusters (PC) and the percentage of noise points returned by DBSCAN. Ground truth has 9 clusters. The top-performing scores are shown in boldface.
Model Clustering Alg. ARI NMI HS CS $S$ PC % Noise
SBERT $k$-Means 0.2210 0.5297 0.7893 0.3986 0.0926
SBERT DBSCAN 0.7462 0.6810 0.7311 0.6374 0.2240 8 14
SBERT HDBSCAN + $k$-NN 0.8593 0.7298 0.6128 0.9023 0.2578 3
SBERT BIRCH 0.2423 0.5608 0.8114 0.4285 0.1037
LayoutLMv1 $k$-Means 0.1533 0.3972 0.5713 0.3044 0.3669
LayoutLMv1 DBSCAN 0.0191 0.1464 0.1393 0.1541 0.5422 5 6.4
LayoutLMv1 HDBSCAN + $k$-NN -0.0567 0.0623 0.0500 0.0827 0.7047 2
LayoutLMv1 BIRCH 0.2403 0.4430 0.6237 0.3435 0.2986
LayoutLMv3 $k$-Means 0.2858 0.5637 0.8136 0.4313 0.2996
LayoutLMv3 DBSCAN 0.7713 0.6995 0.7999 0.6216 0.4493 11 9.6
LayoutLMv3 HDBSCAN + $k$-NN 0.9619 0.8272 0.7695 0.8941 0.5096 4
LayoutLMv3 BIRCH 0.3916 0.5892 0.7902 0.4697 0.3340
DiT $k$-Means 0.0409 0.1796 0.2825 0.1317 0.2011
DiT DBSCAN -0.0386 0.0513 0.0346 0.0992 0.2772 2 11.20
DiT HDBSCAN + $k$-NN 0.0686 0.0599 0.0492 0.0765 0.2841 2
DiT BIRCH 0.0268 0.1535 0.2291 0.1154 0.2276
Donut $k$-Means 0.1883 0.4793 0.7267 0.3576 0.3238
Donut DBSCAN 0.9403 0.8668 0.9269 0.8139 0.2532 11 4
Donut HDBSCAN + $k$-NN 0.9749 0.8906 0.8452 0.9412 0.2739 5
Donut BIRCH 0.1949 0.5046 0.7595 0.3778 0.3324
ColPali $k$-Means 0.2083 0.5300 0.8020 0.3958 0.1215
ColPali DBSCAN 0.6771 0.6218 0.5364 0.7396 0.3440 4 6.40
ColPali HDBSCAN + $k$-NN 0.6646 0.6391 0.5036 0.8741 0.3551 3
ColPali BIRCH 0.4370 0.6656 0.8909 0.5313 0.1370
Gemma3 $k$-Means 0.3164 0.5843 0.8310 0.4505 0.3014
Gemma3 DBSCAN 0.9628 0.8800 0.8738 0.8863 0.4596 6 4.80
Gemma3 HDBSCAN + $k$-NN 0.8780 0.7980 0.8016 0.7944 0.4800 5
Gemma3 BIRCH 0.4067 0.6486 0.8525 0.5234 0.3174
InternVL3 $k$-Means 0.3217 0.4765 0.6629 0.3719 0.4446
InternVL3 DBSCAN 0.5810 0.6267 0.7276 0.5504 0.3846 6 13.20
InternVL3 HDBSCAN + $k$-NN 0.6153 0.6053 0.6343 0.5788 0.3843 5
InternVL3 BIRCH 0.2352 0.5173 0.7688 0.3898 0.3740
Mexican Death Certificates.
@@
Table 5 lists the clustering results in the Mexican death certificate dataset. In the scanned death certificates corpus, the highly repetitive textual content makes the lexical information nearly invariant across forms, restoring the advantage to models that encode spatial and visual structure. The textual clusters produced by SBERT are coarse, with $k$-Means yielding an ARI of 0.221 and a completeness score of 0.3986. Donut attains the best overall performance: HDBSCAN + $k$-NN yields ARI = 0.9749, NMI = 0.8906, homogeneity = 0.8452, and completeness = 0.9412. LayoutLMv3 ranks closely behind, with an ARI of 0.9619 and an NMI of 0.8567. Pure-image DiT, lacking lexical anchoring, collapses to two clusters with 0.039-0.039 ARI and 11% noise, showing that visual tokens alone cannot resolve subtle scanner artefacts. ColPali remains below the other vision-aware encoders (DBSCAN ARI = 0.312, NMI = 0.6771), but still better than LayoutLMv1. DBSCAN performs exceptionally well on Gemma3, achieving an ARI of 0.9628 and an NMI of 0.8800, with 6 clusters and only 4.8% noise. In contrast, InternVL3 delivers only moderate performance.
Implications for Practice and Research.
@@
Methodologically, $k$-Means and BIRCH tend to overstate performance and mask the modality weaknesses that surface as soon as $k$ must be inferred. DiT and Donut lose more than 0.65 ARI in Transformed FATURA when the oracle disappears, while Donut gains 0.78 ARI on Mexican death certificates by switching from $k$-Means to HDBSCAN + $k$-NN. Joint inspection of homogeneity, completeness and the proportion of noise points therefore offers actionable insight that a single aggregate score would obscure, revealing whether an embedding fragments templates or collapses them indistinct, overly broad clusters.
From a practical perspective, the HDBSCAN + $k$-NN hybrid stands out as a strong baseline when labels are unavailable. It consistently matches or exceeds the performance of oracle-$k$ algorithms on two of the three datasets (FATURA and Mexican death certificates), without the constraint of requiring a predefined $k$. Its 0% final noise rate also makes it particularly suitable for production scenarios where every document must be assigned to a cluster. However, its failure on Transformed FATURA highlights a key limitation: no clustering algorithm can overcome a fundamentally misaligned embedding space. In such cases, model choice becomes the dominant factor. SBERT and ColPali retained substantial NMI and ARI under transformation, while DiT, Donut, InternVL3, and even Gemma3 showed significant performance drops. This underscores the importance of embedding robustness to distribution shifts or noise as a critical selection criterion, independent of the choice of clustering algorithm.